OpenStudy (anonymous):

Find a pair of factors for each number by using the difference of two squares. a. 45 b. 77 c. 112 a. b. c.

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

I'm not sure what this question is asking, sorry :/

OpenStudy (anonymous):

ahh ok :/ ty

jigglypuff314 (jigglypuff314):

mmm factor x^2 + 9x + 20 = ?

OpenStudy (anonymous):

im confused :/

jigglypuff314 (jigglypuff314):

to simplify this factor x^2 + 9x + 20 -------------- = x^2 - 25

OpenStudy (anonymous):

i think A

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

yayy :D

jigglypuff314 (jigglypuff314):

mmm if x = √11 foil (6 - x)(6 + x) = ?

OpenStudy (anonymous):

A

jigglypuff314 (jigglypuff314):

close :) you would get 36 - x^2 right? so it would be 36 - (√11)^2 and √11 times √11 = 11 so 36 - 11 = ?

OpenStudy (anonymous):

D. 25 :D

jigglypuff314 (jigglypuff314):

factor out a 2m^6 from (-6m^7 + 20m^6)

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

C

jigglypuff314 (jigglypuff314):

correct! :)

OpenStudy (anonymous):

yaay!! ^-^

jigglypuff314 (jigglypuff314):

mmm the (-3) s can cancel out and denominator can't equal 0

OpenStudy (anonymous):

C :s

jigglypuff314 (jigglypuff314):

mmm once you cancel out the (-3) s your left with q + 2 in the denominator and the denominator can not equal 0

OpenStudy (anonymous):

B :D

jigglypuff314 (jigglypuff314):

mmm q + 2 = 0 ? I'm referring to the "where q =" part

OpenStudy (anonymous):

ohhh makes sense now!

OpenStudy (anonymous):

D

jigglypuff314 (jigglypuff314):

multiply top and bottom by (x-2)(x+2) to get (x+1)(x+1)(x+2) -------------- = (x-1)(x+1) then cancel out a (x+1) from top and bottom

OpenStudy (anonymous):

i think A

jigglypuff314 (jigglypuff314):

Correct! :)

OpenStudy (anonymous):

yay! :-)

jigglypuff314 (jigglypuff314):

multiply top and bottom by (x-8) to get (x-2)(x+2) --------- = (x-2)(x-8) then cancel out the (x-2)

OpenStudy (anonymous):

C

jigglypuff314 (jigglypuff314):

mmm close! when you cancel out the (x-2) s you'll be left with a (x+2) on top right?

OpenStudy (anonymous):

D :)

jigglypuff314 (jigglypuff314):

factor the top and bottom then cancel out a 2

OpenStudy (anonymous):

B

jigglypuff314 (jigglypuff314):

wait... what does it mean by "excluded value" ?

OpenStudy (anonymous):

o.o i have no idea hold on

jigglypuff314 (jigglypuff314):

oh okay then Not B you'll be left with 2(x-2) in denominator denominator cannot equal 0

OpenStudy (anonymous):

D

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

yayy! :D

jigglypuff314 (jigglypuff314):

factor top and bottom cancel out (x-6)

OpenStudy (anonymous):

B :/

jigglypuff314 (jigglypuff314):

when you factor you'll get (x-6)(x+6) --------- -7x (x-6)

OpenStudy (anonymous):

c

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

woo hoo

jigglypuff314 (jigglypuff314):

factor to get (x-6)(x+6) --------- -(x-6)

OpenStudy (anonymous):

hmmm A

jigglypuff314 (jigglypuff314):

Correct! :)

OpenStudy (anonymous):

yay

jigglypuff314 (jigglypuff314):

they are asking what is x when given x^2 + 3x - 10 = 0

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

i think D

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

yayy!! The area of a rectangle is x2 - 2x - 15 and the length of the rectangle is x + 3. Find the width of the rectangle. A. x – 5 B. x2 + 15 C. x + 5 D. (x – 2)(x + 3)

jigglypuff314 (jigglypuff314):

(x-5)(x+3) --------- x+3

OpenStudy (anonymous):

A

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

yayay! :D

jigglypuff314 (jigglypuff314):

idk how to do this one, sorry :/

OpenStudy (anonymous):

how about :/

jigglypuff314 (jigglypuff314):

if x = √6 then what is x + 2x = ?

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

B :s

OpenStudy (anonymous):

im not sure

jigglypuff314 (jigglypuff314):

well what is x + 2x = ?

OpenStudy (anonymous):

3x

jigglypuff314 (jigglypuff314):

then replace x with √6 this works because square roots work like variables :P

OpenStudy (anonymous):

D

jigglypuff314 (jigglypuff314):

mmm 3x then replace x with √6 so 3√6

OpenStudy (anonymous):

ahh ok :/

jigglypuff314 (jigglypuff314):

mmm cross multiply to get x√5 = 2√5 -2 then divide both sides by √5

OpenStudy (anonymous):

C

jigglypuff314 (jigglypuff314):

mmm x = (2√5 -2) / √5

jigglypuff314 (jigglypuff314):

a times a^2 = a^3

OpenStudy (anonymous):

A

jigglypuff314 (jigglypuff314):

mmm? 2a times a^2 = 2a^3

OpenStudy (anonymous):

sorry, B

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

yaa! :))

jigglypuff314 (jigglypuff314):

√39 √6 = √(39*6) = √234 7 times √39 = 7√39

OpenStudy (anonymous):

hmm